OpenStudy (anonymous):

Write a function rule for each table. Days Cost to Rent a Truck 1 34 2 56 3 78 4 100

undeadknight26 (undeadknight26):

for every day it increases by 22...

OpenStudy (anonymous):

but what is the function tho

OpenStudy (anonymous):

c = 22d c = 12d + 22 c = 22d + 22 c = 22d + 12

undeadknight26 (undeadknight26):

i think it is the first one.

OpenStudy (anonymous):

ok ill se

OpenStudy (anonymous):

ok its right thx

undeadknight26 (undeadknight26):

im not sure...

OpenStudy (anonymous):

? Days Cost 2 6 3 9 5 15 6 18 c = d + 3 c = 3d d = 3c c = d + 3d

OpenStudy (anonymous):

idk what is this ????

undeadknight26 (undeadknight26):

hmm i don't understand the table....

OpenStudy (anonymous):

idk how to post the whole table

undeadknight26 (undeadknight26):

not that can you organize it a bit better?

OpenStudy (anonymous):

well gotta goto nextpierod well thx ill hit u up nextpiriod i have ap classses